OpenStudy (anonymous):

find the volume of the solid generated by revolving the line x=-4.The region bounded by the two parabolas x=(y-y^2),x=(y^2-3).

OpenStudy (amistre64):

sounds like a torus...

OpenStudy (amistre64):

need to figure out the solution to the system of equations to know what our "bounds" are gonna be...

OpenStudy (amistre64):

i got an idea, but its just to early in the morning to get it right in me head.....

OpenStudy (amistre64):

the bounds for the integral are gonna be [-1,(3/2)]

OpenStudy (anonymous):

That is an odd region...is it the odd oval shape?

OpenStudy (amistre64):

(S) y-y^2 dy = (1/2)y^2 -(1/3) y^3 |-1-3/2

OpenStudy (amistre64):

its a funky shape fer sure :)

OpenStudy (anonymous):

use circular ring method..

OpenStudy (amistre64):

F(-1) = 5/6

OpenStudy (anonymous):

That shape isn't circular...

OpenStudy (anonymous):

i know the answer but i dont know the solution..

OpenStudy (amistre64):

F(3/2) = 9/8

OpenStudy (anonymous):

v=85.902 cubic units

OpenStudy (amistre64):

your one step ahead of me then :) I am just trying to find the area of the shape

OpenStudy (anonymous):

.plz help me..

OpenStudy (amistre64):

for F(x) + get an area of 7/24....which could be totally wrong, but ill see :)

OpenStudy (anonymous):

I don't understand exactly which region he is asking for...

OpenStudy (amistre64):

G(x) = (S) y^2-3 dy = (1/3)y^3 -3y | -1,(3/2)

OpenStudy (anonymous):

These are the graphs http://www.wolframalpha.com/input/?i=x+%3D+y-y^2%2C+x%3D+y^2-3%2C+x%3D-4

OpenStudy (anonymous):

Why does this site suck at showing links...

OpenStudy (amistre64):

the region for the area is between y=-1 and y=(3/2)

OpenStudy (anonymous):

So it's the odd oval shape in the center http://www.wolframalpha.com/input/?i=x+%3D+y-y^2%2C+x%3D+y^2-3

OpenStudy (amistre64):

i solved the system of equations and got y=-1 and y=(3/2) as solutions

OpenStudy (anonymous):

Yes i see those bounds

OpenStudy (amistre64):

If I did it right; the area of F(x) = 7/24. Now I need to subtract area of G(x) right?

OpenStudy (amistre64):

G(-1) = 8/3

OpenStudy (anonymous):

So where exactly does the line x = -4 come in then

OpenStudy (amistre64):

you spin the area around x=-4 to find the area of the torus shape

OpenStudy (anonymous):

Oh that's the center

OpenStudy (amistre64):

i hope its the center :)

OpenStudy (amistre64):

G(3/2) = -27/8 if Im doing it right :)

OpenStudy (anonymous):

where should i put the strip??is it horizontal or vertical?

OpenStudy (amistre64):

area of G(x) = -145/24 ? Horizontal I beleive

OpenStudy (anonymous):

well it can't be

OpenStudy (amistre64):

total area of shape is F(x) - G(x) = 152/24

OpenStudy (amistre64):

how do we spin it? multiply by circumference of the circle?

OpenStudy (anonymous):

integrate from 0 to 2*pi

OpenStudy (anonymous):

i thought

OpenStudy (amistre64):

I can find the area between curves alright, but after that I get lost :)

OpenStudy (amistre64):

Area = 38/6 = 19/3...thats as far as I can reduce it; if its right

OpenStudy (amistre64):

how do we integrate the spin? Do we need to find the domain of the area are anything?

OpenStudy (anonymous):

I'm not exactly sure

OpenStudy (amistre64):

lol.....comeon..your smarter than me for sure :)

OpenStudy (anonymous):

i thought it was the integral from 0 to 2*pi, but that wouldn't make much sense in this case

OpenStudy (anonymous):

What about using the ring method

OpenStudy (anonymous):

That would probably just be much more difficult

OpenStudy (amistre64):

not that versed in the ring method...

OpenStudy (amistre64):

sounds like we'd integrate with respect to the radius....

OpenStudy (anonymous):

yes the inner radius would be the maximum for the first equation, and the outer radius would be the maximum for the 2nd i believe

OpenStudy (amistre64):

gotta go back and find the bends then :)

OpenStudy (amistre64):

one of those bounds is x=1/4

OpenStudy (amistre64):

is that right? y-y^2 x'=1-2y=0 y=1/2. .5-.5^2 = .25 x=1/4 right?

OpenStudy (anonymous):

My brain hurts lol. Okay so you're deriving y-y^2, solving for y, and plugging it back in?

OpenStudy (amistre64):

y^2 -3 x'=2y=0 y=0 other bound is x=-3

OpenStudy (anonymous):

Yeah that seems right

OpenStudy (amistre64):

yep, only thing I can do :)

OpenStudy (amistre64):

sp we got our bends at x=-3 and x=1/4....now what?

OpenStudy (amistre64):

spin it at x=-4 somehow

OpenStudy (amistre64):

each slice is spin from -3 to 1/4 in dx increments and then add them all up

OpenStudy (anonymous):

Im confused on the area

OpenStudy (amistre64):

make it easier and spin add 4 to x so that we have the middle at 0 and our limits at 1 and the other at 4.25

OpenStudy (anonymous):

Yeah that's what i did actually :)

OpenStudy (amistre64):

which area :) the area between the curves?

OpenStudy (anonymous):

yes, the integral of x = y-y^2 would give the area under the parabola, but how would you remove the sections on the sides to get the torus

OpenStudy (anonymous):

Or i should say...to the right of the parabola

OpenStudy (amistre64):

the two equations meet at the points -1 and 3/2; so these are the bounds we use. to find the area right?

OpenStudy (anonymous):

Ok i see

OpenStudy (amistre64):

back in algebra here :) when y-y^2 = y^2-3 we get solutions to the equations :)

OpenStudy (amistre64):

2y^2 -y -3 =0 (y-...)(y+....)

OpenStudy (amistre64):

I use that to find the area of F(x) and G(x) then: F(x)-G(x) gives us the area we need

OpenStudy (amistre64):

now, since we are talking about spinning a circle; do we (S) 2pi(x)(area)? dx

OpenStudy (amistre64):

it should be the same concept as finding the area between curves; use the lower and upper bounds of x and subtract them?

OpenStudy (amistre64):

cant hurt to try: (S) 2pi(x)(19/3) dx = 19pi(x^2)/3 | 1,4.25

OpenStudy (amistre64):

114.39583333(pi) - 6.3333333333(pi)

OpenStudy (amistre64):

I get 108.0625 pi....but am I right?

OpenStudy (amistre64):

ill need alot more practice for this I guess :)

OpenStudy (amistre64):

v=85.902 cubic units is what the person said to begin with; and that doesnt match mine

OpenStudy (anonymous):

Yeah it doesn't, my brain is too fried for this right now, I've got diff eq in 5 hours, best of luck though

OpenStudy (amistre64):

the key lies in the y axis. distance from the y axis bounds i think..not the x-axis

OpenStudy (anonymous):

what is the integral of 1/((x+2)√(x+3))

OpenStudy (amistre64):

im wrong somewhere :)

OpenStudy (amistre64):

i need more practice on integrals... .

OpenStudy (amistre64):

have you intergrated by parts?

OpenStudy (anonymous):

yeah...but dunno if its correct

OpenStudy (amistre64):

did you do it more than once?

OpenStudy (amistre64):

(S) 1/x+2 dx = ln(x+2) that much works out

OpenStudy (amistre64):

i gotta get some sleep :)....chat with ya later think...

OpenStudy (anonymous):

okay amistre...btw u have taken u=1/x+2?